OpenStudy (cenaida):

A rectangle has an area of 14 square centimeters. Its length is five cm greater than its width. Find the dimensions. A. 2 and 7 B. 4 and 9 C. 6 and 8 D. 7 and 7

OpenStudy (nincompoop):

first determine the area formula of a rectangle

OpenStudy (nincompoop):

A = LW and it says that Area is 14cm, so that makes it 14 = LW the length is 5 cm greater than its width. this means we need to modify length in terms of width 14 = (W+5) W

OpenStudy (cenaida):

so is the with 14 times 5?

OpenStudy (nincompoop):

no

OpenStudy (cenaida):

oh... :/

OpenStudy (cenaida):

then whats the width?

OpenStudy (nincompoop):

2

OpenStudy (nincompoop):

laughing out loud

OpenStudy (cenaida):

ooh..so 2 times 14?

OpenStudy (nincompoop):

14 = (W+5) W distribute then complete the square 6.25 + 14 = W^2 + 5W + 6.25 20.25 = (W+2.5)^2 obtain W by obtaining the square root of both sides of the equation sqrt (20.25) = sqrt[(W + 2.5)^2] 4.5 = W + 2.5 solve for W by adding -2.5 on both sides of the equation 4.5 - 2.5 = W + 2.5 + (-2.5) 2 = W

OpenStudy (cenaida):

so the area is 28?...

OpenStudy (nincompoop):

check: A = L W 14 = (W+5) W 14 = (2+5) 2 14 = (7) 2 14 = 14 true

OpenStudy (nincompoop):

NO. read the problem it begins with "A rectangle has an area of 14 square centimeters. " so the area of the rectangle is 14 cm

OpenStudy (cenaida):

oooh so we need to find the length and width?

OpenStudy (nincompoop):

YEAH!

OpenStudy (nincompoop):

laughing out loud

OpenStudy (cenaida):

LOL ok i kinda get it now then

OpenStudy (cenaida):

so if the width is 2 then whats the lenght?

OpenStudy (nincompoop):

it says the length is 5 cm greater than the width so if width is 2 then length will be 2+5

OpenStudy (cenaida):

oooh 7! so the width is 2 nd the lenght is 7! omg im so dumb ahahaha thnx amor te lo agradesco

OpenStudy (nincompoop):

de nada